# Describing the speed of sound peak of isospin-asymmetric cold strongly interacting matter using effective models

### v1

@@ Abstract
Isospin density, speed of sound, quantum chromodynamics, lattice simulations
I Introduction
@@
In this work we put forward the idea that the peak in the speed of sound can be described by introducing $\mu_{I}$-dependent couplings. That couplings may depend on the properties of the medium is not surprising. In thermo-magnetic QCD it is known that the coupling $\alpha_{s}$ runs with the temperature and field strength [71]. In the context of effective models, this idea has been explored to describe inverse magnetic catalysis using the Nambu–Jona-Lasinio (NJL) model [72, 73, 74, 75, 76] and the Linear Sigma Model with quarks (LSMq) [77, 78]. In the former case, a functional form for the $T$ and $B$-dependent coupling is introduced to mimic asymptotic freedom. In the latter, the $T$ and $B$-dependent couplings are computed from the model itself at one-loop order. In this work we try a different but related approach. We find the $\mu_{I}$ dependence of couplings that describe the isospin density which carries the most direct information of the properties of the speed of sound. In order to respect the thermodynamic identities, we require that the physical pressure does not change when the coupling changes as a function of $\mu_{I}$ [79]. This in turn requires the introduction of a compensating function that implements the pressure-independence under coupling variations. We explicitly explore this scenario using our NJL and LSMq descriptions of isospin imbalanced matter at $T=0$ [80, 81, 82]. The work is organized as follows: In Sec. II we discuss in general the subject of thermodynamic consistency when a parameter that is originally considered as medium-independent is effectively described as medium-dependent. In Secs. III and IV, we introduce the NJL and LSMq models, respectively, and make the couplings become medium-dependent. We fit the isospin density to find the explicit $\mu_{I}$ dependence of the couplings, and apply the thermodynamic consistency procedure to find the description of the rest of the thermodynamic quantities. We show that this procedure provides a good description in particular of the peak structure of the square of the speed of sound as a function of $\mu_{I}$. We finally summarize and discuss our results in Sec. V.
II Thermodynamic consistency
@@
The subject of the thermodynamically consistent description of a physical system, when introducing effective medium dependent quantities, has been addressed in the literature in different contexts [79, 83, 84, 85, 86, 87, 88, 89]. For our purposes, we discuss the subject when introducing a medium-dependent coupling. Under these circumstances, the Hamiltonian, which is originally taken as being medium-independent, now becomes medium-dependent. This leads to an apparent contradiction between the usual thermodynamic relations and their definition using statistical mechanics. To see this, consider the statistical mechanics definitions of the pressure $P$, energy density $\varepsilon$ and number density $n$ for a grand-canonical ensemble described by the inverse temperature $\beta$ and a chemical potential $\mu$
@@
no longer coincides with Eq. (3). The same is also true for other thermodynamic relations such as $\varepsilon$ and the entropy density $s$
$\varepsilon$ $=$ $-P+Ts+\mu n\,,$
$s$ $=$ $\left(\frac{\partial P}{\partial T}\right)_{\mu}\,.$ (5)
A self-consistent solution for this problem is to consider an additional contribution to the Hamiltonian [79], producing an effective Hamiltonian $H_{\textrm{eff}}$. We require that Eqs. (1) – (3) are satisfied by $H_{\textrm{eff}}=H+E^{*}$, where $E^{*}$ also depends on the medium through the model parameter $G(\mu)$, such that $H_{\textrm{eff}}$ does not have an explicit $\mu$-dependence, and thus that
@@
which is equivalent to the constraint in Eq. (II). From this condition and the definition from Eq. (7), we get by integration
@@
III The NJL model at finite isospin chemical potential
@@
To find the physical pressure, we must take into account the considerations of Sec. II. With $B^{*}$ defined in Eq. (10), the total pressure $P$ can be written as
@@
Figure 1: Density dependent coupling $G(\mu_{I})$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$.
Figure 1 presents the behavior found for the isospin chemical potential dependent coupling $G(\mu_{I})$ with the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. The coupling starts as a decreasing function at $\mu_{I}=m_{\pi}$, until reaching a minimum around $\mu_{I}\approx 1.66m_{\pi}$ and increasing thereafter, thus characterizing a non-monotonic behavior. The dashed line indicates the value of $G_{0}$ for the chosen set of parameters.
Figure 2: Scalar condensate $\sigma$ and pion condensate $\Delta$ as functions of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$.
In Fig. 2, we show the chiral condensate $\sigma$ and pion condensate $\Delta$ as functions of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. For comparison both cases, the one with a fixed coupling $G_{0}$ and the one with a density dependent coupling $G(\mu_{I})$, are presented. It can be seen that the values of $G_{0}$ and $G(\mu_{I})$ are the same at the starting value $\mu_{I}=m_{\pi}$ and split thereafter. Both condensates have slightly lower values in the $G=G(\mu_{I})$ case, and the pion condensate $\Delta$ shows some change in its curvature throughout the region of interest.
Figure 3: Normalized isospin density $n_{I}/m_{\pi}^{3}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$.
@@
Figure 4: Speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$.
Finally, in Fig. 4 the square of the speed of sound $c_{s}^{2}$ is shown as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ for both couplings $G_{0}$ and $G(\mu_{I})$, together with the LQCD band for lattice size $32^{3}\times 48$ obtained from Refs. [64, 9]. Again, the fixed coupling case does not reproduce the non-monotonic behavior found in the lattice results. The observed peak is only present in the NJL model if we consider the density dependent coupling $G(\mu_{I})$ obtained by fitting the isospin density.
IV The LSMq at finite isospin chemical potential
@@
$\mathcal{L}$ $=$ $\frac{1}{2}(\partial_{\mu}\sigma)^{2}+\frac{1}{2}(\partial_{\mu}\vec{\pi})^{2}+\frac{a^{2}}{2}(\sigma^{2}+\vec{\pi}^{2})-\frac{\lambda}{4}(\sigma^{2}+\vec{\pi}^{2})^{2}$ (23)
$+$ $i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi-ig\bar{\psi}\gamma^{5}\vec{\tau}\cdot\vec{\pi}\psi-g\bar{\psi}\psi\sigma,$
@@
$\pi_{-}=\frac{1}{\sqrt{2}}(\pi_{1}+i\pi_{2}),\quad\pi_{+}=\frac{1}{\sqrt{2}}(\pi_{1}-i\pi_{2}).$ (27)
@@
$\partial_{\mu}\to D_{\mu}=\partial_{\mu}+i\mu_{I}\delta_{\mu}^{0},\quad\partial^{\mu}\to D^{\mu}=\partial^{\mu}-i\mu_{I}\delta_{0}^{\mu},$ (28)
@@
$V_{\rm tree}=-\frac{a^{2}}{2}\left(v^{2}+\Delta^{2}\right)+\frac{\lambda}{4}\left(v^{2}+\Delta^{2}\right)^{2}-\frac{1}{2}\mu_{I}^{2}\Delta^{2}-hv.$ (37)
@@
$E_{\Delta}^{u}$ $=$ $\left\{\left(\sqrt{k^{2}+m_{f}^{2}}+\mu_{I}\right)^{2}+g^{2}\Delta^{2}\right\}^{1/2},$ (39a)
$E_{\Delta}^{d}$ $=$ $\left\{\left(\sqrt{k^{2}+m_{f}^{2}}-\mu_{I}\right)^{2}+g^{2}\Delta^{2}\right\}^{1/2},$ (39b)
@@
Notice tha this choice differs from the one in Eq. (12). Equation (38) is ultraviolet divergent. To identify the divergent terms, we work in the approximation whereby the fermion energies, Eqs. (IV), are expanded in powers of $\mu_{I}^{2}$
@@
$+$ $\frac{\mu_{I}^{2}g^{2}\Delta^{2}}{(k^{2}+m_{f}^{2}+g^{2}\Delta^{2})^{3/2}}\Big{)}$
@@
$\sum_{f=u,d}V_{f}^{1}$ $=$ $2N_{c}\frac{g^{4}\left(v^{2}+\Delta^{2}\right)^{2}}{(4\pi)^{2}}\left[\frac{1}{\epsilon}+\frac{3}{2}+\ln\left(\frac{\Lambda^{2}/g^{2}}{v^{2}+\Delta^{2}}\right)\right]$ (42)
-- $2N_{c}\frac{g^{2}\mu_{I}^{2}\Delta^{2}}{(4\pi)^{2}}\left[\frac{1}{\epsilon}+\ln\left(\frac{\Lambda^{2}/g^{2}}{v^{2}+\Delta^{2}}\right)\right],$
@@
$\!\!\!\!\!\!\!\!\left.\frac{\partial V_{\rm tree}}{\partial v}=\left[\lambda v^{3}-(a^{2}-\lambda\Delta^{2})v-h\right]\right|_{v_{0},\,\Delta_{0}}$ $=$ 00 0 (43a)
$\!\!\!\!\!\!\!\!\left.\frac{\partial V_{\rm tree}}{\partial\Delta}=\left[\lambda\Delta^{2}-(\mu_{I}^{2}-\lambda v^{2}+a^{2})\right]\right|_{v_{0},\,\Delta_{0}}$ $=$ $0$ (43b)
Notice that the second of Eqs. (IV) admits a real, non-vanishing solution, only when
@@
which means that a non-zero isospin condensate is developed only when, for positive values of the isospin chemical potential, the latter is larger than the vacuum pion mass. This is what we identify as the condensed phase. The simultaneous solutions of Eqs. (IV) are
@@
Hereafter, we refer to the expressions in Eq. (IV) as the classical solution.
Figure 5: Density dependent coupling $g(\mu_{I})$ of the LSMq as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. For comparison we also show the case of a constant coupling $g$ with the dashed line.
@@
Figure 6: Normalized isospin density $n_{I}/m_{\pi}^{3}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of LSMq fixed coupling $g$ and density dependent coupling $g(\mu_{I})$.
@@
$\lambda=\frac{m_{\sigma}^{2}-m_{\pi}^{2}}{2f_{\pi}^{2}}.$
@@
For $m_{q}=270$ MeV, $m_{\sigma}=450$ MeV, $m_{\pi}=131.7$ MeV and $f_{\pi}=92.4$ MeV, one readily obtains $\lambda=10.84$, $a=274.29$ MeV and $g=2.9$. These values are chosen to provide a reasonably good description of the LQCD data, particularly for the pressure [9, 64]. In what follows, we concentrate on finding the $\mu_{I}$ induced correction to $g$ treating $\lambda$ as not being influenced by the medium. The reason for this choice is than in the approximation where fermions induce loop effects whereas bosons are treated as average fields, the coupling that involves fermions, namely $g$, should be the only coupling that receives medium modifications.
Figure 5 shows the isospin chemical potential dependent effective coupling $g(\mu_{I})$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ obtained by fitting the LSMq expression for the isospin density to the LQCD data from Ref. [64, 9]. Notice that, as opposed to the result for the NJL model, $g(\mu)$ is a monotonically increasing function of $\mu_{I}$. Figure 6 shows the LSMq normalized isospin density as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the obtained $g(\mu_{I})$ from the LQCD data from Ref. [64, 9]. For comparison, we also show the LSMq result using a constant coupling $g$.
@@
Figure 7 shows the LSMq square of the speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the density dependent coupling $g(\mu_{I})$, compared to the LQCD data from Ref. [64, 9]. For comparison, we also show the LSMq result using a constant coupling $g$. Notice that the peak for $\mu_{I}\sim 1.5\ m_{\pi}$ is well reproduced by employing the found profile of $g(\mu_{I})$.
@@
Figure 7: Square of the speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the LSMq density dependent coupling $g(\mu_{I})$. For comparison, shown is also $c_{s}^{2}$ using the constant value for $g$.
V Summary and discussion

### v2

@@ Abstract
keywords:
Isospin density, speed of sound, quantum chromodynamics, lattice simulations, effective models
1 Introduction
@@
In this work we put forward the idea that the peak in the speed of sound can be described by introducing $\mu_{I}$-dependent couplings. That couplings may depend on the properties of the medium is not surprising. In thermo-magnetic QCD it is known that the coupling $\alpha_{s}$ runs with the temperature and field strength [71]. In the context of effective models, this idea has been explored to describe inverse magnetic catalysis using the Nambu–Jona-Lasinio (NJL) model [72, 73, 74, 75, 76] and the Linear Sigma Model with quarks (LSMq) [77, 78]. In the former case, a functional form for the $T$ and $B$-dependent coupling is introduced to mimic asymptotic freedom. In the latter, the $T$ and $B$-dependent couplings are computed from the model itself at one-loop order. In this work we try a different but related approach. We find the $\mu_{I}$ dependence of couplings that describe the isospin density which carries the most direct information of the properties of the speed of sound, utilizing as inputs the lattice results from Refs. [9, 64]. However, the introduction of a medium dependent parameter to the system via the coupling has consequences for the thermodynamic identities. In this particular case, the isospin density, as computed by its statistical mechanical definition, would not coincide with that computed from the thermodynamic relations. Under the assumption that these calculations should coincide, a compensating function can be introduced to the system’s Hamiltonian (and consequently to the effective pressure) and be subsequently determined by identifying the coupling variations. The subject of the thermodynamic consistency of systems with medium dependent parameters has been explored in the literature to a considerable extent. In Ref. [79], the authors generalized the main ideas within a study of a phenomenological model with a gluon medium dependent mass. Further applications in a wide range of contexts may be found in Refs. [80, 81, 82, 83, 84, 85, 86, 87]. Here, we explicitly explore this scenario using our two-flavor NJL and LSMq descriptions of isospin imbalanced matter at $T=0$ [88, 89, 90]. Notice that the LQCD results, used in this work to compare with our model calculations, are performed using 2+1 flavors. In the LQCD calculation, the strange quark is included at the action level with zero chemical potential ($\mu_{s}=0$). While the operator $n_{I}$ depends only on $\mu_{u}$ and $\mu_{d}$, the expectation value of the isospin density, $\left\langle n_{I}\right\rangle$, also depends on $\mu_{s}$ via the determinant weights in the path integral. This means that in a perturbative calculation, the effect of the strange quark on $\left\langle n_{I}\right\rangle$ can only be captured at higher loop orders involving $s$-quark bubbles. In the spirit of a perturbative calculation, these higher order contributions are thus suppressed with respect to a one-loop order calculation, such as what we present in this work. Therefore a comparison between our two-flavor model analysis and the referenced LQCD data is justified within physical grounds since our model calculations capture the leading order effect. Moreover, it has been observed in Refs. [36, 89] that the two-flavor NJL model and chiral perturbation theory calculations are in better agreement with LQCD results in comparison with their respective three-flavor versions.
@@
The work is organized as follows: In Sec. 2 we discuss in general the subject of thermodynamic consistency when a parameter that is originally considered as medium-independent is effectively described as medium-dependent. In Secs. 3 and 4, we introduce the NJL and LSMq models, respectively, and make the couplings become medium-dependent. We fit the isospin density to find the explicit $\mu_{I}$ dependence of the couplings, and apply the thermodynamic consistency procedure to find the description of the rest of the thermodynamic quantities. We show that this procedure provides a good description in particular of the peak structure of the square of the speed of sound as a function of $\mu_{I}$. We finally summarize and discuss our results in Sec. 5.
2 Thermodynamic consistency
@@
The subject of the thermodynamically consistent description of a physical system, when introducing effective medium dependent quantities, has been addressed in the literature in different contexts [79, 80, 81, 82, 83, 84, 85, 86, 87]. For our purposes, we discuss the subject when introducing a medium-dependent coupling. Under these circumstances, the Hamiltonian, which is originally taken as being medium-independent, now becomes medium-dependent. This leads to an apparent contradiction between the usual thermodynamic relations and their definition using statistical mechanics. To see this, consider the statistical mechanics definitions of the pressure $P$, energy density $\varepsilon$ and number density $n$ for a grand-canonical ensemble described by the inverse temperature $\beta$ and a chemical potential $\mu$
@@
no longer coincides with Eq. (3). This apparent ambiguity in the definition of the number density may then have consequences for the energy density $\varepsilon$
$\varepsilon$ $=$ $-P+Ts+\mu n\,,$ (5)
where $s=\left(\partial P/\partial T\right)_{\mu}$ is the entropy density. A self-consistent solution for this problem is to consider an additional contribution to the Hamiltonian [79], producing an effective Hamiltonian $H_{\textrm{eff}}$. We require that Eqs. (1) – (3) are satisfied by $H_{\textrm{eff}}=H+E^{*}$, where $E^{*}$ also depends on the medium through the model parameter $G(\mu)$, such that $H_{\textrm{eff}}$ does not have an explicit $\mu$-dependence, and thus that
@@
which is equivalent to the constraint in Eq. (2). From this condition and the definition from Eq. (7), we get by integration
@@
3 The NJL model at finite isospin chemical potential
@@
To find the physical pressure, we must take into account the considerations of Sec. 2. With $B^{*}$ defined in Eq. (10), the total pressure $P$ can be written as
@@
Figure 1: Density dependent coupling $G(\mu_{I})$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$
Figure 1 presents the behavior found for the isospin chemical potential dependent coupling $G(\mu_{I})$ with the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. The coupling starts as a decreasing function at $\mu_{I}=m_{\pi}$, until reaching a minimum around $\mu_{I}\approx 1.66m_{\pi}$ and increasing thereafter, thus characterizing a non-monotonic behavior. The dashed line indicates the value of $G_{0}$ for the chosen set of parameters. It can be seen that the values of $G_{0}$ and $G(\mu_{I})$ are the same at the starting value $\mu_{I}=m_{\pi}$ and split thereafter.
Figure 2: Scalar condensate $\sigma$ and pion condensate $\Delta$ as functions of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$
In Fig. 2, we show the chiral condensate $\sigma$ and pion condensate $\Delta$ as functions of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. For comparison both cases, the one with a fixed coupling $G_{0}$ and the one with a density dependent coupling $G(\mu_{I})$, are presented. Both condensates have slightly lower values in the $G=G(\mu_{I})$ case, and the pion condensate $\Delta$ shows some change in its curvature throughout the region of interest.
Figure 3: Normalized isospin density $n_{I}/m_{\pi}^{3}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$
@@
Figure 4: Speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of fixed coupling $G_{0}$ and density dependent coupling $G(\mu_{I})$
Finally, in Fig. 4 the square of the speed of sound $c_{s}^{2}$ is shown as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ for both couplings $G_{0}$ and $G(\mu_{I})$, together with the LQCD band for lattice size $32^{3}\times 48$ obtained from Refs. [64, 9]. Again, the fixed coupling case does not reproduce the non-monotonic behavior found in the lattice results. The observed peak is only present in the NJL model if we consider the density dependent coupling $G(\mu_{I})$ obtained by fitting the isospin density. Notice that the predicted behavior of the coupling can be traced back to that of the isospin density used for the fitting procedure. From Fig. 3, we observe that the NJL model with a fixed coupling already provides a good description of the LQCD results at low $\mu_{I}$. The model curve with constant $G$ separates from the LQCD curve from above at intermediate values of $\mu_{I}$ to then come closer again to the LQCD curve for large values of $\mu_{I}$. This is reflected on the found properties of the coupling $G(\mu_{I})$ in Fig. 1, as it decreases from the nominal value $G_{0}$ for low values of $\mu_{I}$, reaches a minimum around the inflection point observed in $n_{I}$ at intermediate values of $\mu_{I}$, and strengthens back again as the isospin density grows closer to the constant coupling case at high $\mu_{I}$. We also emphasize that the numerical changes in $G$ do not deviate too much from the nominal value $G_{0}$. Its maximum value is $\sim$ 4.8 GeV-2 and its minimum is $\sim$ 4.5 GeV-2 throughout the considered region, which corresponds to an overall variation of about $\sim 6\%$. Nonetheless, this dependence of $G$ with $\mu_{I}$ is directly linked to the non-monotonic behavior of the speed of sound, as the same good description found for $n_{I}$ reflects in the pressure and energy density due to thermodynamic consistency.
4 The LSMq at finite isospin chemical potential
@@
$\mathcal{L}$ $=$ $\frac{1}{2}(\partial_{\mu}\sigma)^{2}+\frac{1}{2}(\partial_{\mu}\vec{\pi})^{2}+\frac{a^{2}}{2}(\sigma^{2}+\vec{\pi}^{2})$ (23)
-- $\frac{\lambda}{4}(\sigma^{2}+\vec{\pi}^{2})^{2}+i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi-ig\bar{\psi}\gamma^{5}\vec{\tau}\cdot\vec{\pi}\psi$
-- $g\bar{\psi}\psi\sigma,$
@@
$\pi_{\pm}=\frac{1}{\sqrt{2}}(\pi_{1}\mp i\pi_{2}).$ (25)
@@
$\partial_{\mu}\to D_{\mu}=\partial_{\mu}+i\mu_{I}\delta_{\mu}^{0},$
$\partial^{\mu}\to D^{\mu}=\partial^{\mu}-i\mu_{I}\delta_{0}^{\mu}.$ (26)
@@
$V_{\rm tree}$ $=$ $-\frac{a^{2}}{2}\left(v^{2}+\Delta^{2}\right)+\frac{\lambda}{4}\left(v^{2}+\Delta^{2}\right)^{2}$ (35)
-- $\frac{1}{2}\mu_{I}^{2}\Delta^{2}-hv.$
@@
$E_{\Delta}^{u}$ $=$ $\left\{\left(\sqrt{k^{2}+m_{f}^{2}}+\mu_{I}\right)^{2}+g^{2}\Delta^{2}\right\}^{1/2}\!\!\!,$
$E_{\Delta}^{d}$ $=$ $\left\{\left(\sqrt{k^{2}+m_{f}^{2}}-\mu_{I}\right)^{2}+g^{2}\Delta^{2}\right\}^{1/2}\!\!\!,$ (37)
@@
Notice that this choice differs from the one in Eq. (12). Equation (36) is ultraviolet divergent. To identify the divergent terms, we work in the approximation whereby the fermion energies, Eqs. (37), are expanded in powers of $\mu_{I}^{2}$
@@
$+$ $\frac{\mu_{I}^{2}g^{2}\Delta^{2}}{(k^{2}+m_{f}^{2}+g^{2}\Delta^{2})^{3/2}}\Big{)}.$
@@
$\sum_{f=u,d}V_{f}^{1}$ $=$ $2N_{c}\frac{g^{4}\left(v^{2}+\Delta^{2}\right)^{2}}{(4\pi)^{2}}\left[\frac{1}{\epsilon}+\frac{3}{2}\right.$ (40)
$+$ $\left.\ln\left(\frac{\Lambda^{2}/g^{2}}{v^{2}+\Delta^{2}}\right)\right]-2N_{c}\frac{g^{2}\mu_{I}^{2}\Delta^{2}}{(4\pi)^{2}}\left[\frac{1}{\epsilon}\right.$
$+$ $\left.\ln\left(\frac{\Lambda^{2}/g^{2}}{v^{2}+\Delta^{2}}\right)\right],$
@@
$\left.\frac{\partial V_{\rm tree}}{\partial v}=\left[\lambda v^{3}-(a^{2}-\lambda\Delta^{2})v-h\right]\right|_{v_{0},\,\Delta_{0}}$ $=$ $0,$
$\left.\frac{\partial V_{\rm tree}}{\partial\Delta}=\Delta\left[\lambda\Delta^{2}-(\mu_{I}^{2}-\lambda v^{2}+a^{2})\right]\right|_{v_{0},\,\Delta_{0}}$ $=$ $0$
Notice that the second of Eqs. (41) admits a real, non-vanishing solution, only when
@@
which means that a non-zero isospin condensate is developed only when, for positive values of the isospin chemical potential, the latter is larger than the vacuum pion mass. This is what we identify as the condensed phase. The simultaneous solutions of Eqs. (41) are
@@
Hereafter, we refer to the expressions in Eq. (43) as the classical solution.
Figure 5: Density dependent coupling $g(\mu_{I})$ of the LSMq as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$. For comparison we also show the case of a constant coupling $g$ with the dashed line
@@
Figure 6: Normalized isospin density $n_{I}/m_{\pi}^{3}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$, considering the cases of LSMq fixed coupling $g$ and density dependent coupling $g(\mu_{I})$
@@
$\lambda=\frac{m_{\sigma}^{2}-m_{\pi}^{2}}{2f_{\pi}^{2}}.$ (50)
@@
For $m_{q}=270$ MeV, $m_{\sigma}=450$ MeV, $m_{\pi}=131.7$ MeV and $f_{\pi}=92.4$ MeV, one readily obtains $\lambda=10.84$, $a=274.29$ MeV and $g=2.9$. These values are chosen to provide a reasonably good description of the LQCD data, particularly for the pressure [9, 64]. In what follows, we concentrate on finding the $\mu_{I}$ induced correction to $g$ treating $\lambda$ as not being influenced by the medium. The reason for this choice is that in our approach, fermions induce loop effects whereas bosons are treated as average fields. Thus, the coupling that involves fermions, namely $g$, should be treated as the only coupling receiving medium modifications.
@@
Figure 5 shows the isospin chemical potential dependent effective coupling $g(\mu_{I})$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ obtained by fitting the LSMq expression for the isospin density to the LQCD data from Ref. [64, 9]. Figure 6 shows the LSMq normalized isospin density as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the obtained $g(\mu_{I})$ from the LQCD data from Ref. [64, 9]. For comparison, we also show the LSMq result using a constant coupling $g$. Figure 7 shows the LSMq square of the speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the density dependent coupling $g(\mu_{I})$, compared to the LQCD data from Ref. [64, 9]. For comparison, we also show the LSMq result using a constant coupling $g$. Notice that the peak for $\mu_{I}\sim 1.5\ m_{\pi}$ is well reproduced by employing the found profile of $g(\mu_{I})$.
We notice from Fig. 5 that, as opposed to the result for the NJL model, $g(\mu_{I})$ is a monotonically increasing function of $\mu_{I}$. The figure also shows that the constant value (represented by the dashed line) corresponds to the average of $g(\mu_{I})$ in the considered range for the analysis. This behavior can be understood recalling that a constant value can only describe the gross features of the isospin density dependence on $\mu_{I}$. In order to describe the detailed features, it can then be expected that for some range of $\mu_{I}$ the $g$-values are below this average whereas for another range the $g$-values are above, so that on average $g$ takes its given constant value. Here, $g(\mu_{I})$ starts below the constant value because it is computed from the properties of the isospin density, which in turn is computed as the derivative of the pressure with respect to $\mu_{I}$. According to the results of Ref. [90], the latter starts off from $\mu_{I}=m_{\pi}$ with a smaller value of the derivative as compared with the LQCD data. It may also be noted that the value of $g(\mu_{I}/m_{\pi}=1)$ does not coincide with that of the NJL model coupling $G(\mu_{I}/m_{\pi}=1)$, first because the two constants do not have the same units - $g$ is dimensionless whereas $G$ has units of inverse mass squared. It is nevertheless important to point out that if the coupling $G$ is multiplied by the square of the ultraviolet cutoff $\Lambda$, the other dimensionful quantity associated to vacuum properties, one obtains $G\Lambda^{2}\simeq 2.07$, which coincides with the value of $g(\mu_{I}/m_{\pi}=1)$. From a more detailed viewpoint, $g$ and $G$ are quantities that capture the physics of systems possessing chiral symmetry in a different manner: In the NJL model, as can be seen from Eq. (17), $G$ (or rather, its inverse) represents the strength with which the condensates contribute to the systems energy. In the LSMq, $g$ has a similar role, as can be seen from Eq. (40), but this behavior is obtained using in addition the explicit expression for the quark mass $m_{q}$ in terms of the coupling $g$ and the chiral condensate $v$.
@@
Figure 7: Square of the speed of sound $c_{s}^{2}$ as a function of the normalized isospin chemical potential $\mu_{I}/m_{\pi}$ using the LSMq density dependent coupling $g(\mu_{I})$. For comparison, shown is also $c_{s}^{2}$ using the constant value for $g$
5 Summary and discussion